### v1

# Federated nnU-Net for Privacy-Preserving Medical Image Segmentation

Abstract
The nnU-Net framework has played a crucial role in medical image segmentation and has become the gold standard in multitudes of applications targeting different diseases, organs, and modalities. However, so far it has been used primarily in a centralized approach where the data collected from hospitals are stored in one center and used to train the nnU-Net. This centralized approach has various limitations, such as leakage of sensitive patient information and violation of patient privacy. Federated learning is one of the approaches to train a segmentation model in a decentralized manner that helps preserve patient privacy. In this paper, we propose FednnU-Net, a federated learning extension of nnU-Net. We introduce two novel federated learning methods to the nnU-Net framework – Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) – and experimentally show their consistent performance for breast, cardiac and fetal segmentation using 6 datasets representing samples from 18 institutions. Additionally, to further promote research and deployment of decentralized training in privacy-constrained institutions, we make our plug-n-play framework public. The source-code is available at https://github.com/faildeny/FednnUNet.
data privacy, federated learning, image segmentation, nnU-Net, nnUNet
1.  Introduction
Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5, 6]. Deep learning methods, especially convolutional neural networks (CNNs), have shown great success in this field. One of the most effective approaches is the no-new-U-Net framework (nnU-Net) [7] which, over the years, has established itself as the go-to framework for medical image segmentation task, evidenced by its multiple first-place finishes on public leaderboards and its widespread adoption as the foundation for further model development across diverse segmentation challenges [8, 9]. nnU-Net extends the widely adopted U-Net architecture [10], which utilizes an encoder-decoder structure with skip connections to capture both semantic and localization information. Going beyond the original U-Net, nnU-Net automatically configures network architectures tailored to a given dataset, provides a robust training schedule and data augmentation strategy, as well as postprocessing methods to further improve the final performance. This “no-new” principle eliminates much of the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [7, 11, 12, 13]. Despite these advancements, the development of robust segmentation models often requires aggregating extensive datasets from multiple institutions to capture diverse patient populations and imaging protocols. However, the consolidation of such data is frequently impeded by stringent data privacy regulations, ethical considerations, and logistical challenges associated with data sharing. Federated learning (FL) [14] has emerged as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data. Federated learning, however, presents significant challenges when data are heterogeneous across institutions. Traditional FL approaches typically assume that all participants train identical models on similarly distributed data, yet in real-world scenarios, clients often have different computational resources and dataset characteristics. To address this disparity, methods such as HeteroFL [15] have been introduced to enable federated training among clients with varying model capacities, while more recent efforts [16] have focused on mitigating label noise and inconsistencies inherent in federated settings. Collectively, these approaches underscore the growing need for robust FL techniques that can simultaneously manage model and data heterogeneity. Regardless of the significant federated developments in medical domain [17, 18, 19, 20, 21, 22, 23], those methods have not been explicitly tailored for medical image segmentation tasks using nnU-Net. Moreover, nnU-Net’s plug-and-play design, which has made it the gold standard in centralized settings, no longer functions seamlessly in federated environments, necessitating tailored adaptations to overcome the challenges of decentralized, heterogeneous data. To bridge the gap between the decentralized learning paradigms and the robust capabilities of nnU-Net framework, we introduce FednnU-Net, federated extension of nnU-Net that enables collaborative training while preserving data privacy. To the best of our knowledge, it is the first fully-federated and privacy-preserving implementation of nnU-net. We propose two novel methods–Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg). These methods address the challenges posed by variations in dataset properties across different centers, ensuring that each institution maximizes the benefit from shared knowledge training. Our experiments demonstrate that both methods consistently outperform local, non-collaborative training setups for 2D and 3D segmentation tasks across six public and non-public datasets. By extending nnU-Net’s automated pipeline to federated learning and introducing methods to handle data heterogeneity, this work aims to advance the development of robust, scalable, and privacy-preserving medical image segmentation solutions.
2.  Methods
We propose two novel methods to federate nnU-Net’s pipeline for image segmentation: FFE and AsymFedAvg. These two approaches address the challenges of decentralized, heterogeneous data in model training while maintaining nnU-Net’s self-configuring advantages.
2.1.  nnU-Net vs FednnU-Net
nnU-Net is a multi-stage segmentation pipeline that begins with a preparation stage, where dataset fingerprint extraction is performed and a customized training plan (including a tailored network architecture) is generated. This is followed by a training stage and concludes with a post-training stage that selects the best-performing models and optimal image post-processing techniques. To extend nnU-Net into a federated learning setting, we integrate our proposed methods at key points in this pipeline. The first method, FFE, is incorporated into the preparation stage to ensure that data characteristics from all participating institutions are captured and considered when generating the training plan. The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous architectures. The following sections provide detailed descriptions of each method. For clarity, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process, while ”server” describes the coordinating agent responsible for communication and aggregation of inputs from the nodes.
2.2.  Federated Fingerprint Extraction
Figure 1: The Federated Fingerprint Extraction (FFE) process. Each node computes a local fingerprint from its dataset and shares it with the server, which then aggregates them into a global fingerprint. Finally, the server redistributes this global fingerprint to all nodes for local nnU-Net configuration. This process happens once at the beginning of the training stage.
The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup. This approach effectively approximates the configuration behaviour in a centralized environment. The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies. The FFE approach comprises two key steps:
  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net. This local fingerprint encloses the local dataset’s pixel intensity-based metrics, as well as spatial characteristics (such as spacing and voxel size) (see Table 1).
  • Global Fingerprint Aggregation:
    The nodes share their local fingerprint with the server, where these get aggregated to form a global fingerprint. Explicitly, the characteristics are aggregated as follows:
    $\hat{F}={\hat{f}k=g_{k}({f_{k}^{i}}{i=1}^{N})|k\in\mathcal{K}}$ (1)
    where $\hat{F}$ denotes the aggregated fingerprint, $f_{k}^{i}$ represents the $k-th$ fingerprint key from node $i$, and $g_{k}$ is the corresponding aggregation function. Table 1 lists the corresponding $g_{k}$ functions for each fingerprint key. The weighted_mean is calculated with weights proportional to data samples quantity from each participating node.
Table 1: Dataset fingerprint characteristics and aggregation Functions.
Fingerprint Key Aggregation Function $g_{k}$
Pixel intensity
max global max
min global min
mean weighted_mean
median weighted_mean
std weighted_mean
percentile_00_5 weighted_mean
percentile_99_5 weighted_mean
Shape
median_relative_size_after_cropping weighted_mean
shapes_after_crop concatenate
spacings concatenate
Fig. 1 showcases an overview of the FFE process pipeline. As a final step, each node calculates its own nnU-Net training plan based on the received global fingerprint and local hardware characteristics. This process ensures that all nodes will be optimizing either highly compatible or, in most cases, identical network architectures (if the default nnU-Net’s memory constraint has not been altered between nodes).
2.3.  Asymmetric Federated Averaging
In scenarios where nnU-Net is deployed across federated nodes, variations in model architecture can arise due to differences in data characteristics or computational capabilities. To address this inherent heterogeneity, we extend the classical FedAvg approach by developing an Asymmetric Federated Averaging strategy (AsymFedAvg). This method enables nodes to share only the parts of the model that have a common space with the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the locally extracted dataset fingerprints of the centers. A set of example configurations obtained for each data center is presented in Table 2.
Table 2: Default nnU-Net training configurations for each center in MAMA-MIA dataset
Parameter ISPY1 ISPY2 DUKE NACT
2D Configuration
Batch Size 49 32 12 49
Patch Size [256,256] [320,320] [512,512] [256,256]
Spacing [0.78,0.78] [0.68,0.68] [0.70,0.70] [0.70,0.70]
U-Net Stages 7 7 8 7
Features/Stage [32-512] [32-512] [32-512] [32-512]
Let $\mathcal{K}={1,\dots,K}$ be the set of participating nodes in the federation. For each node $k\in\mathcal{K}$ we define a local model state dictionary as:
$S^{k}={(l_{i}^{k},\theta_{i}^{k})|i\in\mathcal{L}^{k}}$ (2)
where $l_{i}^{k}$ represents the layer identifier, $\theta_{i}^{k}$ denotes the corresponding parameters for node $k$, and $\mathcal{L}^{k}$ is the set of layers in node $k$’s model.
We define a layer matching function $\mathcal{C}$ which identifies the common layers of the models participating in the federation as:
$\small\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}$ (3)
For each compatible layer $l\in\mathcal{C}({S^{k}}_{k=1}^{K})$, the aggregated parameters $\hat{\theta}_{l}$ are computed as:
$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k}$ (4)
where $\mathcal{K}_{l}$ is the set of nodes that contain layers $l$.
The aggregated model state dictionary $\hat{S}$ is then constructed as:
$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}$ (5)
During each federation round $t$, the update process goes as follows:
Algorithm 1 Asymmetric Model Aggregation
1:Each node $k$ sends its state dictionary $S^{k}_{t}$ to the server
2:The server identifies compatible layers using the matching function $\mathcal{C}$
3:The server computes $\hat{S}_{t}$ using the averaging procedure
4:for each node $k$ do
5:     for each layer $l$ in model do
6:         if $l\in\mathcal{C}(\{S^{k}\}_{k=1}^{K})$ then
7:              $\theta_{l,t+1}^{k}\leftarrow\hat{\theta}_{l,t}$
8:         else
9:              $\theta_{l,t+1}^{k}\leftarrow\theta_{l,t}^{k}$
10:         end if
11:     end for
12:end for
The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space (compatible layers), handling the diverse nature of medical imaging data across different centers. Fig. 2 depicts an overview of the AsymFedAvg pipeline and the process of updating the heterogeneous architectures in each training round.
Figure 2: The Asymmetric Federated Averaging (AsymFedAvg) Pipeline process. The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline. During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg. Finaly, at the end of each training round, the server distributes the aggregated layers of the network back into the nodes. This process happens throughout the whole training.
2.4.  Implementation
2.4.1 Federated Learning Environment
For the federated communication, we select the Flower framework [24], which is an open-source library designed to streamline federated learning tasks on a cluster of machines. Flower’s high-level API eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it already supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [25] and Federated SGD (FedSGD) [14].
Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, as it allows seamless integration of local models while maintaining centralized coordination on the server.
2.4.2 FednnU-Net
Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real and simulated environments. Its modular structure enables seamless integration not only with the original nnU-Net but also with its customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup.
2.4.3 Distributed Setup
Our experimental framework consists of 4 computational nodes with varying hardware and system specifications (Table 3), resembling real-world clinical scenarios in low-resource countries. In particular, the availability of GPU device with 8GB of memory, is the requirement for each node in order to run the default nnU-Net training setup. The proposed framework was deployed and tested independently across all nodes, although the final model training is done on a single, multi-GPU cluster to simplify experimental control.
Table 3: Hardware specifications of the computational nodes.
Node CPU (Cores/GHz)* GPU (Memory)** OS ***
Node 1 i7-8700K (12/3.20) GTX 1080 (8GB) Ubuntu 22.04
Node 2 i7-8700K (12/3.20) GTX 1080 (8GB) Windows 10
Node 3 i7-9700K (8/3.60) RTX 2080S (8GB) Ubuntu 22.04
Node 4 i7-9700 (8/3.00) RTX 2080S (8GB) Ubuntu 22.04
*GPU: Graphics Processing Unit, **CPU: Central Processing Unit, ***OS: Operating system
3.  Datasets
We evaluate our proposed methods on three distinct clinical applications, using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts. The following sections provide a detailed description of the datasets employed in each use case.
(a) MAMA-MIA
(b) M&M
(c) AIMIX
Figure 3: Example images with annotations from each data center used in the study.
3.1.  Breast MRI
3.1.1 MAMA-MIA
The MAMA-MIA dataset [26] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking AI algorithms in breast cancer imaging. It comprises 1,506 cases acquired between 1995 and 2016 from four cohorts within The Cancer Imaging Archive (TCIA): ISPY-1 (n=171), ISPY-2 (n=980), Duke (n=291), and NACT (n=64). The annotation process focused on delineating primary tumors and non-mass-enhanced lesions within a single breast, using the first post-contrast phase as the reference—a phase typically associated with maximum enhancement of cancerous tissues that offers superior contrast between lesions and surrounding healthy tissue compared to later phases. However, significant imaging differences among these cohorts (Table 4) pose challenges for training robust models. For example, the NACT cohort predominantly features scans with a 2 mm slice thickness and an average of 60 slices per case, whereas Duke’s samples are mostly acquired with 1 mm thickness and include an average of 169 slices per case. Moreover, the centers used different scanner models and field strengths, further contributing to the variability in imaging protocols.
Table 4: Summary of MRI acquisition parameters across the four centers (ISPY1, ISPY2, DUKE, and NACT) in the MAMA-MIA breast cancer dataset [26].
Center Vendor Scanner Model
Slice
Thickness (mm)
Number of
Slices
Field
Strength (T)
Number of
Studies
ISPY1
GE (67.3%)
Siemens (25.7%)
Philips (7.0%)
Signa Genesis (60.2%)
Other (39.8%)
2.4 [1.5, 4.0] 64 [44, 256]
1.5 (100.0%)
3.0 (0.0%)
171
ISPY2
GE (62.3%)
Siemens (25.7%)
Philips (11.9%)
Signa HDxt (54.7%)
Avanto (12.6%)
Other (32.7%)
2.0 [0.8, 3.0] 106 [52, 256]
1.5 (73.0%)
3.0 (27.0%)
980
DUKE
GE (60.5%)
Siemens (39.5%)
Other (55.7%)
Avanto (24.1%)
Signa HDxt (20.3%)
1.1 [1.0, 2.5] 169 [60, 256]
1.5 (46.7%)
3.0 (53.3%)
291
NACT GE (100%) Signa Genesis (100%) 2.0 [2.0, 2.4] 60 [46, 64]
1.5 (100.0%)
3.0 (0.0%)
64
3.1.2 EuCanImage:
Through the European project EuCanImage [^1] https://eucanimage.eu/, we access a non-public breast MRI collection comprising 574 samples from three different countries. The EuCanImage dataset represents a multi-institutional effort to create standardized imaging biobanks for cancer research and serves here as an external test set to validate our methodology on data acquired under protocols that differ markedly from those in MAMA-MIA. Detailed specifications of the acquisition parameters are provided in Table 5. Notably, while MAMA-MIA includes both axial and sagittal acquisitions, the EuCanImage collection consists exclusively of axial slices. In addition, images in EuCanImage are predominantly acquired from scanner models that differ from those represented in MAMA-MIA, and a significantly larger fraction of the scans are captured at 3T field strength. Although the spatial image characteristics are largely similar between the two datasets, these differences in scanner models, field strengths, and imaging orientations lead to distinct scan appearances, potentially challenging the robustness of segmentation algorithms.
Table 5: Summary of MRI acquisition parameters across the three centers (GUMED, KAUNO, and HCB) in the EUCanImage breast MRI dataset.
Center Vendor Scanner Model
Slice
Thickness (mm)
Pixel
Spacing (mm)
Field
Strength (T)
Number of
Studies
GUMED (Poland) Siemens Aera (100%) 1.00–1.49 1.00–1.49 1.5 (100%) 30
KAUNO (Lithuania)
Philips (88.4%)
Siemens (11.6%)
Ingenia (88.4%)
Avanto (11.6%)
1.50–1.99 (84.5%) 0.50–0.99 (96.6%)
1.5 (11.6%)
3.0 (88.4%)
232
HCB (Spain)
GE (69.2%)
Siemens (30.8%)
Signa HDxt (63.8%)
Aera (22.8%)
Others (13.4%)
2.00–2.49 (87.2%) 0.50–0.99 (98.1%)
1.5 (90.7%)
3.0 (9.3%)
312
3.2.  Cardiac MRI
3.2.1 M&Ms
To further validate our approach, we leverage the Multi-center, Multi-vendor & Multi-disease (M&Ms) cardiac MRI datasets. As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [27];[28] by excluding the overlapping samples based on anonymized patient IDs provided by each institution. This curation allows to reach the full potential of this data collection without the risk of replicated samples. The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct institutions. We make this collection publicly available [^2] This data resource will be released upon the acceptance of this article.. The M&Ms datasets are particularly valuable for validation due to their inherent heterogeneity (Table 6), they include images from four different vendors (Siemens, GE, Philips, Canon), various scanner models (from older systems like Symphony to newer ones like Vantage Orian), and different field strengths (1.5T and 3.0T). This multi-center, multi-vendor nature mirrors the real-world variability in clinical settings, making these datasets ideal for testing the generalizability of our approach.
The technical heterogeneity across centers is evident in variations of scanner models, field strengths and different acquisition protocols. This diversity presents similar challenges to those encountered in our MAMA-MIA analysis, particularly in terms of standardization and harmonization of imaging parameters.
Table 6: Average specifications of images acquired in different centers across the merged collection of M&Ms 1 [27] and M&Ms 2 [28] challenge datasets.
Center Vendor Scanner Model
Slice
Thickness (mm)
Number of
Slices
Field
Strength (T)
Number of
Studies
Hospital Vall d’Hebron Siemens Symphony TIM (63.18%) Avanto Fit (15.48%) Other (21.34%) 9.5 13 1.5 (97.91%) 3.0 (2.09%) 239
Clinica Sagrada Familia Philips Achieva 10 13 1.5 98
Universitätsklinikum Hamburg-Eppendorf Philips Achieva 10 11 1.5 51
Hospital Universitari Dexeus GE Signa Excite (75.24%) Signa HDxt (23.81%) Signa Explorer (0.95%) 9.9 12 1.5 (88.57%) 3.0 (11.43%) 105
Hospital Sant Pau Canon Vantage Orian 10 13 1.5 50
3.2.2 ACDC:
The Automated Cardiac Diagnosis Challenge (ACDC) dataset consists of 150 cardiac short-axis MRI examinations acquired at the University Hospital of Dijon (France) [3]. The dataset is evenly distributed across five groups (30 cases each): normal subjects (NOR), patients with myocardial infarction (MINF), dilated cardiomyopathy (DCM), hypertrophic cardiomyopathy (HCM), and abnormal right ventricle (RV). This dataset is particularly valuable for validation due to its well-defined pathological diversity and standardized acquisition protocols. It is used for external validation of models trained on M&Ms dataset.
3.3.  Fetal Ultrasound
3.3.1 AIMIX
With a focus on resource-constrained environments, the AIMIX [^3] https://aimix-erc.eu/ dataset was developed to advance inclusive image analysis in prenatal care. This dataset addresses critical challenges, including the scarcity of African medical data, limited computational resources, and a lack of clinical expertise. The dataset comprises pregnancy blind-sweeps 2D ultrasound scans, along with associated patient metadata such as age and ethnicity. The data span multiple international sites, including Algeria, Egypt, Malawi and Spain, reflecting diverse geographic and healthcare settings.
In this study, we focus on abdominal segmentation, a crucial step in prenatal ultrasound analysis. To enable the evaluation on this multi-center dataset, we provide a set of 130 standard-plane images with masks corresponding to the fetal abdomen [^4] This data resource will be released upon the acceptance of this article.. Our data processing workflow involved detecting the optimal abdominal plane from the each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [29], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician. Table 7 lists the number of samples and the mean resolution for each center.
Table 7: Summary of the AIMIX and ACOUSLIC-AI datasets, including number of samples, mean resolution, nnU-Net patch size, and UNet stages for each dataset.
Center
Mean Resolution
(h x w)
Number of
Studies
Algeria 418x500 25
Egypt 450x400 20
Malawi 400x500 25
Spain 1228x1276 60
SierraLeone (External) 561x743 178
3.3.2 ACOUSLIC-AI Sierra Leone:
For the external validation, we utilize the publicly available samples from ACOUSLIC-AI [^5] https://acouslic-ai.grand-challenge.org/datasets/ Challenge dataset [30]. We use a collection of 178 cases collected in Sierra Leone from three Public Health Units (PHUs). With a resolution of 744×562 pixels, the images from Sierra Leone sit between the low-end devices used in Algeria and Malawi and the high-end scanners employed in Spain. This intermediate imaging quality makes the dataset an ideal candidate for assessing model generalizability across varying image qualities and clinical environments, ensuring robust validation of AI performance.
4.  Experiments
4.1.  Experimental setups
The training process is conducted across four different setups:
  • Single-center Training:  Each node independently trains its own nnU-Net model using only its local dataset. This setup serves as a baseline, reflecting the performance of a node/institution without participation in the federation.
  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized model. This setup represents the theoretical upper bound of performance, though it is typically infeasible in real-world scenarios due to privacy constraints.
  • Federated Fingerprint Training:   Utilizing our FFE approach (2), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node. As a result, each training site obtains identical network architecture during the nnU-Net’s self-configuration process. This approach approximates the nnU-Net’s behavior in the centralized training setup.
  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, allowing nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
For our experimentation, we maintain nnU-Net’s default configuration parameters as described in [7]. We test the 2D U-Net architecture on breast MRI and fetal ultrasound data and the 3D architecture on a cardiac MRI use case, consistently employing the default training length of 1000 epochs and using the last model checkpoint for evaluation across all federated nodes. Following other recent works on medical image segmentation [7, 31, 32, 33], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95). For internal validation, we report results as the average of a 5-fold cross-validation, while for external testing, we follow the original nnU-Net approach by generating ensemble predictions from all five folds. Moreover, each experiment uses identical data splits—for example, the n-th fold in the centralized setup corresponds directly to the set of n-th folds from centers used in the local setup. By adhering to these protocols, we try to minimize the influence of other factors on the final results, ensuring that any observed performance variations are primarily due to our federated learning adaptations rather than modifications to the base framework.
4.2.  Results
Table 8: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
Method ISPY1 ISPY2 DUKE NACT external GUMED external KAUNO external HCB
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
Centralized 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
Table 9: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
Method Hospital Vall d’Hebron Clinica Sagrada Familia Universitätsklinikum Hamburg-Eppendorf Hospital Dexeus Hospital Sant Pau external ACDC
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
Centralized 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
Table 10: Quantitative results for multi-center AIMIX and ACOUSLIC-AI datasets. HD95 values reported in pixels due to missing spacing information. Best results for privacy-preserving methods are in bold.
Method Algeria Egypt Malawi Spain external Sierra Leone
Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
Single-center 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
Centralized 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
(a) MAMA-MIA cross-center validation matrix
(b) M&Ms cross-center validation matrix
(c) AIMIX cross-center validation matrix
Figure 4: Cross-center DSC scores for isolated model training predictions across multiple centers.
4.2.1 Breast MRI
Table 8 presents the segmentation performance for the breast lesion MAMA-MIA dataset. Both proposed methods, FFE and AsymFedAvg, outperform the local training of individual centers in terms of DSC and HD95 across all datasets. Specifically, FFE method achieves an average DSC improvement of 3.2% on ISPY1, 0.3% on ISPY2, 3.7% on DUKE, and a notable 10.9% on NACT. Additionally, FFE maintains DSC scores close to the centralized approach, while surpassing it in terms of HD95, effectively bridging the performance gap in a decentralized framework. Interestingly for DUKE, AsymFedAvg outperforms every other method in terms of HD95. Since the DUKE dataset comprises samples of substantially different spatial characteristics (pixel resolution approximately 2 times higher, see Table 4), the adaptive nature of the asymmetric aggregation approach allows this center to maintain its initial architecture, while benefiting from the other centers. In this case, the HD95 gets reduced in contrast to unified architecture methods (FFE and centralized setting). The cross-center validation matrix in Fig. 4 further exposes the dataset’s heterogeneity, where both ISPY2 and NACT trained models struggle to generalize well for the DUKE dataset, with the only exception of ISPY2, which has image resolution much closer to DUKE. Sample predictions presented in Fig. 5 show the strong improvement of segmentation performance for NACT dataset, which struggles to perform well with its locally trained nnU-Net model. For DUKE, HD95 improvement can be observed for AsymFedAvg model prediction. External validation results for the EuCanImage dataset show that federated training with the FFE method achieves DSC scores on par with the centralized model. While the segmentation performance for the KAUNO center is comparable to the results achieved within MAMA-MIA dataset, demonstrating good generalizability. The GUMED center exhibits a notable decrease in performance, which is most likely attributed to discrepancies in annotation protocols rather than the data itself. An example of such difference is illustrated in Fig. 7, where fine-grained model predictions fail to align with coarse ground truth regions provided by experts.
Figure 5: Sample predictions of each method for Breast Cancer datasets.
4.2.2 Cardiac MRI
For M&Ms, similarly to previous dataset, both proposed federated learning methods outperform the local, non-collaborative training setup across both tested metrics. The only exception being Clinica Sagrada Familia, for which the HD95 is lower in the local setup by a minimal margin. FFE method consistently achieves the highest scores in terms of DSC for both internal and external datasets, while the adaptive AsymFedAvg does not reflect substantial benefits from federated training. Most likely, it is caused by dominantly homogeneous nature of M&Ms dataset, in which all data centers share similar sample characteristics making a common architecture method more suitable. This effect can be observed as well in the cross-center evaluation of locally trained models (see Fig. 4) where each center managed to provide acceptable results for samples from other, unseen institutions. Moreover, it is important to note that the FFE method, despite the additional privacy requirement manages to provide almost identical performance to the centralized training scenario. Sample predictions for cardiac MRI dataset can be observed in Fig. 6. All tested methods provide consistent results for central slices while having more visible disparities in regions close to the heart’s apex (sample predictions from ACDC) and heart’s atria (sample predictions from Vall d’Hebron).
Figure 6: Sample predictions of each method for cardiac MRI datasets.
4.2.3 Fetal Ultrasound
For fetal abdomen segmentation task, considerable differences can be observed between FFE and AsymFedAvg methods; see Table 10. Due to strong data heterogeneity, i.e. discrepancies in capture devices and image resolution between centers (Table 7), the models trained with data from Algeria, Egypt or Malawi do not generalize well for higher resolution samples from Spain; see Fig. 4c. In this case, having a unified training strategy for all centers does not prove to be the optimal solution. Allowing the largest center, i.e. Spain, to keep its original architecture allows to get a more appropriate model to the characteristics of the external center, i.e. Sierra Leone, while still benefiting from the federated training and contribution of other centers. This effect can be observed on sample predictions presented in Fig. 7. The single-center nnU-Net model often struggled to identify the right region for abdomen delineation, which can be especially observed in the external validation samples from Sierra Leone.
Figure 7: Sample predictions of each method for AIMIX dataset.
5.  Discussion
Our experiments across three diverse imaging use cases - breast MRI (MAMA-MIA), cardiac MRI (M&Ms), and fetal ultrasound (AIMIX) - show that FednnU-Net can reliably match, or even surpass conventional centralized strategies. These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while respecting the complexities of real-world clinical collaborations, such as data privacy and varied institution-specific protocols. Moreover, our findings underscore the critical need for methods that effectively handle data heterogeneity during federated model training.
Whilst other, federated state-of-the-art segmentation methods may achieve higher accuracy with highly customized solutions for specific tasks, the primary objective of this work is to provide an easy-to-use, privacy-preserving solution accessible to researchers with varying levels of expertise. In doing so, we prioritize the broader applicability and real-world viability of the framework over achieving the absolute highest segmentation performance.
The increasing emphasis on privacy-preserving, decentralized settings is rapidly becoming the standard in modern AI solutions — especially within healthcare, where patient data confidentiality is paramount. FednnU-Net aligns with this trend by enabling collaborative training without compromising data privacy. However, several aspects require further investigation:
  1. 1.
    Scalability and Client Diversity: Future studies should test FednnU-Net with a larger number of nodes to better mimic real-world deployments, where numerous and geographically dispersed institutions may challenge model convergence and communication.
  2. 2.
    Heterogeneous Computational Resources: Given that clients often have different processing capabilities, it is crucial to evaluate FednnU-Net under varied hardware constraints. Although our framework already supports such a scenario, this dimension was not explored in this current study.
  3. 3.
    Enhanced Privacy Mechanisms: Beyond our current privacy measures, integrating techniques such as differential privacy [34, 35] and secure multiparty computation (SMPC) could further reduce the risk of data leakage, even in aggregated updates.
In summary, although more accurate medical segmentation techniques exist, FednnU-Net fills a vital gap by balancing segmentation performance with ease-of-use and privacy preservation. As federated learning continues to mature and becomes an increasingly standard practice in healthcare, addressing these limitations through further evaluation and enhancements will be essential for broader adoption and improved robustness in clinical applications.
6.  Conclusion
In this study, we present FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that bridges the gap between privacy-sensitive setups and high-performance model training. Our framework introduces two methods: Federated Fingerprint Extraction (FFE) and Asymetric Federated Averaging (AsymFedAvg) that unlock the potential of distributed model training under privacy constraints. FFE enables construction of one, universal model architecture akin to centralized training, ensuring robustness across data centers. Alternatively, AsymFedAvg allows tailored model customization for each center to better align with local data characteristics. Our results demonstrate that both FFE and AsymFedAvg offer distinct performance benefits depending on the multi-center dataset characteristics, either achieving performance close to the centralized setting or surpassing it, particularly in highly diverse datasets. On top of that, both approaches preserve nnU-Net’s hallmark self-configuring capabilities, and due to our modular open-source implementation, they allow researchers to seamlessly translate their nnU-Net-derived centralized segmentation methods to the federated setting. Furthermore, we open source our framework and release data enhancing two public dataset collections to allow reproducibility of results. We hope that our contributions provide a robust foundation for further development and evaluation of medical image segmentation models in privacy-constrained settings.

### v2

@@
 The nnU-Net framework has played a crucial role in medical image segmentation and has become the gold standard in multitudes of applications targeting different diseases, organs, and modalities.
-However, so far it has been used primarily in a centralized approach where the data collected from hospitals are stored in one center and used to train the nnU-Net.
-This centralized approach has various limitations, such as leakage of sensitive patient information and violation of patient privacy.
-Federated learning is one of the approaches to train a segmentation model in a decentralized manner that helps preserve patient privacy.
-In this paper, we propose FednnU-Net, a federated learning extension of nnU-Net.
-We introduce two novel federated learning methods to the nnU-Net framework – Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) – and experimentally show their consistent performance for breast, cardiac and fetal segmentation using 6 datasets representing samples from 18 institutions.
-Additionally, to further promote research and deployment of decentralized training in privacy-constrained institutions, we make our plug-n-play framework public.
-The source-code is available at https://github.com/faildeny/FednnUNet.
+However, so far it has been used primarily in a centralized approach where the collected data is stored in the same location where nnU-Net is trained.
+This centralized approach has various limitations, such as potential leakage of sensitive patient information and violation of patient privacy.
+Federated learning has emerged as a key approach for training segmentation models in a decentralized manner, enabling collaborative development while prioritising patient privacy.
+In this paper, we propose FednnU-Net, a plug-and-play, federated learning extension of the nnU-Net framework.
+To this end, we contribute two federated methodologies to unlock decentralized training of nnU-Net, namely, Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg).
+We conduct a comprehensive set of experiments demonstrating high and consistent performance of our methods for breast, cardiac and fetal segmentation based on a multi-modal collection of 6 datasets representing samples from 18 different institutions.
+To democratize research as well as real-world deployments of decentralized training in clinical centres, we publicly share our framework at https://github.com/faildeny/FednnUNet.
 data privacy, federated learning, image segmentation, nnU-Net, nnUNet
-1.  Introduction
-Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5, 6].
-Deep learning methods, especially convolutional neural networks (CNNs), have shown great success in this field.
+Introduction
+Medical image segmentation plays a key role in clinical workflows to accurately delineate anatomical structures and help with diagnosis, treatment planning, and disease monitoring [1, 2, 3, 4, 5].
+Deep learning methods, especially convolutional neural networks (CNNs) [6, 7], have shown great success in this field.
 One of the most effective approaches is the no-new-U-Net framework (nnU-Net) [8] which, over the years, has established itself as the go-to framework for medical image segmentation task, evidenced by its multiple first-place finishes on public leaderboards and its widespread adoption as the foundation for further model development across diverse segmentation challenges [9, 10].
@@
 Going beyond the original U-Net, nnU-Net automatically configures network architectures tailored to a given dataset, provides a robust training schedule and data augmentation strategy, as well as postprocessing methods to further improve the final performance.
-This “no-new” principle eliminates much of the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [7, 11, 12, 13].
+This “no-new” (nn) principle largely eliminates the manual effort associated with hyperparameter tuning and architectural engineering, rendering nnU-Net both powerful and highly convenient across a broad range of biomedical tasks [8, 12, 13, 14].
 Despite these advancements, the development of robust segmentation models often requires aggregating extensive datasets from multiple institutions to capture diverse patient populations and imaging protocols.
 However, the consolidation of such data is frequently impeded by stringent data privacy regulations, ethical considerations, and logistical challenges associated with data sharing.
-Federated learning (FL) [14] has emerged as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data.
-Federated learning, however, presents significant challenges when data are heterogeneous across institutions.
-Traditional FL approaches typically assume that all participants train identical models on similarly distributed data, yet in real-world scenarios, clients often have different computational resources and dataset characteristics.
-To address this disparity, methods such as HeteroFL [15] have been introduced to enable federated training among clients with varying model capacities, while more recent efforts [16] have focused on mitigating label noise and inconsistencies inherent in federated settings.
-Collectively, these approaches underscore the growing need for robust FL techniques that can simultaneously manage model and data heterogeneity.
-Regardless of the significant federated developments in medical domain [17, 18, 19, 20, 21, 22, 23], those methods have not been explicitly tailored for medical image segmentation tasks using nnU-Net.
-Moreover, nnU-Net’s plug-and-play design, which has made it the gold standard in centralized settings, no longer functions seamlessly in federated environments, necessitating tailored adaptations to overcome the challenges of decentralized, heterogeneous data.
-To bridge the gap between the decentralized learning paradigms and the robust capabilities of nnU-Net framework, we introduce FednnU-Net, federated extension of nnU-Net that enables collaborative training while preserving data privacy.
-To the best of our knowledge, it is the first fully-federated and privacy-preserving implementation of nnU-net.
-We propose two novel methods–Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg).
+Federated learning (FL) [15] was introduced as a promising solution to these challenges by facilitating collaborative model training across multiple institutions without the need to exchange raw data.
+Although nnU-Net has become a standard for medical image segmentation in centralized settings, its use in FL is challenging due to its dataset-specific customization which involves preprocessing, normalization, model choice, and training schedules based on the underlying data [8].
+The model architecture itself includes design decisions regarding the size and stride of convolutional kernels, as well as the number of network blocks.
+Such level of customization fundamentally conflicts with standard FL methods like FedAvg [15] that require uniform model architectures and training procedures across clients.
+To solve these limitations, we introduce FednnU-Net–the first fully federated privacy-preserving implementation of nnU-Net.
+FednnU-Net is specifically designed to maintain the automated and out-of-the-box nature of nnU-Net while overcoming the challenges associated with decentralized, heterogeneous data.
+To this end, we propose two methods—Federated Fingerprint Extraction (FFE) and Asymmetric Federated Averaging (AsymFedAvg) to unlock distributed training of nnU-Net.
 These methods address the challenges posed by variations in dataset properties across different centers, ensuring that each institution maximizes the benefit from shared knowledge training.
 Our experiments demonstrate that both methods consistently outperform local, non-collaborative training setups for 2D and 3D segmentation tasks across six public and non-public datasets.
-By extending nnU-Net’s automated pipeline to federated learning and introducing methods to handle data heterogeneity, this work aims to advance the development of robust, scalable, and privacy-preserving medical image segmentation solutions.
-2.  Methods
-We propose two novel methods to federate nnU-Net’s pipeline for image segmentation: FFE and AsymFedAvg.
-These two approaches address the challenges of decentralized, heterogeneous data in model training while maintaining nnU-Net’s self-configuring advantages.
-2.1.  nnU-Net vs FednnU-Net
-nnU-Net is a multi-stage segmentation pipeline that begins with a preparation stage, where dataset fingerprint extraction is performed and a customized training plan (including a tailored network architecture) is generated.
-This is followed by a training stage and concludes with a post-training stage that selects the best-performing models and optimal image post-processing techniques.
-To extend nnU-Net into a federated learning setting, we integrate our proposed methods at key points in this pipeline.
-The first method, FFE, is incorporated into the preparation stage to ensure that data characteristics from all participating institutions are captured and considered when generating the training plan.
-The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous architectures.
-The following sections provide detailed descriptions of each method.
-For clarity, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process, while ”server” describes the coordinating agent responsible for communication and aggregation of inputs from the nodes.
-2.2.  Federated Fingerprint Extraction
+By extending nnU-Net’s automated pipeline to FL and introducing methods to handle data heterogeneity, this work advances the development of robust, scalable, and privacy-preserving medical image segmentation solutions.
+Furthermore, its modular design also unlocks access to bringing a wide range of nnU-Net extensions to the federated settings, including improved training efficiency through e.g.
+curriculum learning [16], enhanced model explainability [17], continual learning [18], or uncertainty estimation [19, 20, 21].
+Related Work
+Federated learning in medical imaging
+FL has been widely adopted in medical imaging for tasks such as disease classification, lesion detection, and organ segmentation across multiple institutions [22, 23, 24].
+A range of federated segmentation methods have been introduced to tackle key challenges such as data heterogeneity, communication inefficiency, and limited generalization across sites.
+Notable examples include FedEvi [25], which employs evidence-based aggregation to account for uncertainty in model updates; FedGS [26], which utilizes gradient statistics to enhance robustness; and Fed-MENU [27], which integrates model ensembling with uncertainty-aware learning to improve stability.
+Transformer-based [28] methods leverage self-attention mechanisms and present high performance in distributed training for positron emission tomography (PET) and computed tomography (CT) segmentation [29].
+Additionally, recent approaches like FedFMS [30] incorporate foundation models like SAM [31] and MedSAM [32] into FL through fine-tuning, aiming to harness large-scale pretrained representations for downstream segmentation tasks in a privacy-preserving context.
+Despite the overall progress, these methods present limitations in terms of practical applicability.
+Many rely on highly customized architectures, complex aggregation strategies, or specialized pretraining, which can pose significant barriers for researchers without profound expertise in machine learning.
+In contrast to nnU-Net framework, Transformer- and foundational model-based approaches offer much less flexibility in terms of computational resources, making them less feasible in typical clinical environments with constrained hardware.
+These constraints limit their out-of-the-box usability and, as a result, chances of clinical application.
+There remains a strong need for universal FL methods that are robust across datasets, computationally efficient, and readily applicable by a broad range of users—including those with limited resources or domain-specific requirements.
+Heterogeneous model aggregation
+Conventional FL assumes identical models and similarly distributed data across clients.
+However, real-world deployments often involve heterogeneous datasets and varying compute capacities.
+To address this, HeteroFL [33] allows training subnetworks of different widths from a shared supernetwork.
+While effective for standard CNNs, it is incompatible with nnU-Net, which adapts model depth, resolution, and preprocessing to dataset-specific features.
+FjORD [34] introduces Ordered Dropout for training nested submodels, enabling dynamic scaling.
+Yet, it relies on a fixed architecture, limiting its applicability to nnU-Net’s dataset-driven design.
+FedRolex [35] improves fairness by cyclically training submodels from a global network.
+However, it still assumes all client models are subsets of a shared template—an assumption that nnU-Net’s self-configuring framework violates.
+Since none of the existing methods accommodates the flexible and adaptive nature of nnU-Net, we develop a new federated approach tailored specifically to nnU-Net’s architecture and requirements.
+Methods
+We propose two tailored methods to federate nnU-Net’s pipeline for image segmentation, namely, FFE and AsymFedAvg.
+These two approaches are tailored towards addressing the challenges of decentralized, heterogeneous data in model training while simultaneously allowing to maintain nnU-Net’s self-configuring advantages.
+nnU-Net vs FednnU-Net
+nnU-Net is a multi-stage segmentation pipeline initiated by a preparation stage, where dataset fingerprint extraction is performed and a customized training plan is generated, which includes a tailored segmentation network architecture.
+This is followed by a training stage before concluding with a post-training stage that selects the best-performing models alongside optimal image post-processing techniques.
+To extend nnU-Net into a FL framework, we integrate our proposed methods in the respectively corresponding components of this pipeline.
+The first method, FFE, is incorporated into the preparation stage, ensuring that data characteristics from all participating institutions are captured and considered when generating the training plan.
+The second method, AsymFedAvg, operates during the training stage and facilitates the aggregation of model weight updates across nodes with heterogeneous network architectures.
+With the following sections providing detailed descriptions of each method, we use the terms ”node” and ”client” interchangeably to refer to each institution participating in the decentralized training process.
+We further denote the term ”server” as the coordinating agent responsible for communication and aggregation of inputs from the nodes.
+Federated Fingerprint Extraction
+The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup.
+This approach effectively approximates the configuration behavior in a centralized environment.
+The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies.
+The FFE approach comprises two key steps:
+  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net.
+This local fingerprint encloses the local dataset’s spatial characteristics such as spacing and voxel size.
+  • Global Fingerprint Aggregation: The nodes share their local fingerprint with the server, where these get aggregated to form a global federated fingerprint z.
+Specifically, new lists of shapes_after_crop and spacings are created by the concatenation of their local representations.
 Figure 1: The Federated Fingerprint Extraction (FFE) process.
@@
 This process happens once at the beginning of the training stage.
-The proposed FFE method enables the analysis of the federated datasets and the configuration of nnU-Net, allowing to determine a unified, one-fits-all training strategy for all data centers in the federated setup.
-This approach effectively approximates the configuration behaviour in a centralized environment.
-The dataset’s fingerprint is a crucial component of the nnU-Net’s adaptive configuration, summarizing the key dataset characteristics that are used to determine the model configuration and preprocessing strategies.
-The FFE approach comprises two key steps:
-  • Local Fingerprint Extraction: Each participating node generates their dataset fingerprint by invoking the extraction process of nnU-Net.
-This local fingerprint encloses the local dataset’s pixel intensity-based metrics, as well as spatial characteristics (such as spacing and voxel size) (see Table 1).
-  • Global Fingerprint Aggregation:
-    The nodes share their local fingerprint with the server, where these get aggregated to form a global fingerprint. Explicitly, the characteristics are aggregated as follows:
-    $\hat{F}={\hat{f}k=g_{k}({f_{k}^{i}}{i=1}^{N})|k\in\mathcal{K}}$ (1)
-    where $\hat{F}$ denotes the aggregated fingerprint, $f_{k}^{i}$ represents the $k-th$ fingerprint key from node $i$, and $g_{k}$ is the corresponding aggregation function.
-Table 1 lists the corresponding $g_{k}$ functions for each fingerprint key.
-The weighted_mean is calculated with weights proportional to data samples quantity from each participating node.
-Table 1: Dataset fingerprint characteristics and aggregation Functions.
-Fingerprint Key Aggregation Function $g_{k}$
-Pixel intensity
-max global max
-min global min
-mean weighted_mean
-median weighted_mean
-std weighted_mean
-percentile_00_5 weighted_mean
-percentile_99_5 weighted_mean
-Shape
-median_relative_size_after_cropping weighted_mean
-shapes_after_crop concatenate
-spacings concatenate
 Fig.
 1 showcases an overview of the FFE process pipeline.
-As a final step, each node calculates its own nnU-Net training plan based on the received global fingerprint and local hardware characteristics.
-This process ensures that all nodes will be optimizing either highly compatible or, in most cases, identical network architectures (if the default nnU-Net’s memory constraint has not been altered between nodes).
-2.3.  Asymmetric Federated Averaging
+As a final step, each node calculates its own nnU-Net training plan based on the combination of the received global fingerprint and the local hardware characteristics.
+This process ensures that all nodes will be optimizing either highly compatible or, in the case of nodes with sufficiently similar local memory and hardware constraints, identical network architectures.
+Asymmetric Federated Averaging
 In scenarios where nnU-Net is deployed across federated nodes, variations in model architecture can arise due to differences in data characteristics or computational capabilities.
-To address this inherent heterogeneity, we extend the classical FedAvg approach by developing an Asymmetric Federated Averaging strategy (AsymFedAvg).
-This method enables nodes to share only the parts of the model that have a common space with the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
-At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the locally extracted dataset fingerprints of the centers.
-A set of example configurations obtained for each data center is presented in Table 2.
+To address this inherent heterogeneity, we extend the classical FedAvg approach by developing a novel Asymmetric Federated Averaging strategy (AsymFedAvg).
+In contrast to other heterogeneous aggregation methods mentioned in Heterogeneous model aggregation, our method allows the aggregation of models with different lengths and layer configurations.
+AsymFedAvg enables nodes to share only the parts of the model that are common across the rest of the federation, effectively eliminating the need to constrain all nodes to a single, identical architecture.
+At each participating node, we initialize nnU-Net’s adaptive configuration pipeline to generate node-specific U-Net architectures based on the aforementioned locally extracted dataset fingerprints of the centers.
+For illustration, a set of example configurations obtained from each data center is presented in Table 1.
+Figure 2: Overview of the Asymmetric Federated Averaging (AsymFedAvg) Training Pipeline.
+The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline.
+During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg.
+At the end of each training round, the server distributes the aggregated layers of the network back into the nodes.
+This process is respectively repeated throughout the entire model training process.
 Table 1: Default nnU-Net training configurations for each center in MAMA-MIA dataset
@@
 Let $\mathcal{K}={1,\dots,K}$ be the set of participating nodes in the federation. For each node $k\in\mathcal{K}$ we define a local model state dictionary as:
-$S^{k}={(l_{i}^{k},\theta_{i}^{k})|i\in\mathcal{L}^{k}}$ (2)
+$S^{k}={(l_{i}^{k},\theta_{i}^{k})|i\in\mathcal{L}^{k}},$ (1)
 where $l_{i}^{k}$ represents the layer identifier, $\theta_{i}^{k}$ denotes the corresponding parameters for node $k$, and $\mathcal{L}^{k}$ is the set of layers in node $k$’s model.
 We define a layer matching function $\mathcal{C}$ which identifies the common layers of the models participating in the federation as:
-$\small\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}$ (3)
+$\mathcal{C}({S^{k}}_{k=1}^{K})={l|\forall k,j\in\mathcal{K}:l\in\mathcal{L}^{k}\cap\mathcal{L}^{j}\land\text{dim}(\theta_{l}^{k})=\text{dim}(\theta_{l}^{j})}.$ (2)
 For each compatible layer $l\in\mathcal{C}({S^{k}}_{k=1}^{K})$, the aggregated parameters $\hat{\theta}_{l}$ are computed as:
-$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k}$ (4)
+$\hat{\theta}_{l}=\frac{1}{|\mathcal{K}l|}\sum{k\in\mathcal{K}_{l}}\theta_{l}^{k},$ (3)
 where $\mathcal{K}_{l}$ is the set of nodes that contain layers $l$.
 The aggregated model state dictionary $\hat{S}$ is then constructed as:
-$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}$ (5)
-During each federation round $t$, the update process goes as follows:
+$\hat{S}={(l,\hat{\theta}l)|l\in\mathcal{C}({S^{k}}{k=1}^{K})}.$ (4)
+During each federation round $t$, the update process goes as presented in Algorithm  1.
 Algorithm 1 Asymmetric Model Aggregation
@@
 12:end for
-The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space (compatible layers), handling the diverse nature of medical imaging data across different centers.
+Notation:
+  • $S^{k}_{t}$: State dictionary of node $k$ at round $t$
+  • $\mathcal{C}$: Layer compatibility function
+  • $\hat{\theta}_{l,t}$: Aggregated parameters for layer $l$
+The AsymFedAvg allows nodes to maintain their unique architectural components while sharing knowledge via their common space of compatible layers, handling the diverse nature of medical imaging data across different centers.
 Fig.
-2 depicts an overview of the AsymFedAvg pipeline and the process of updating the heterogeneous architectures in each training round.
-Figure 2: The Asymmetric Federated Averaging (AsymFedAvg) Pipeline process.
-The nodes of the federated setup generate a local nnU-Net from the nnU-Net pipeline.
-During training, the server receives the local nnU-Net weights of each node and aggregates them utilizing the Aggregation Function of AsymFedAvg.
-Finaly, at the end of each training round, the server distributes the aggregated layers of the network back into the nodes.
-This process happens throughout the whole training.
-2.4.  Implementation
-2.4.1 Federated Learning Environment
-For the federated communication, we select the Flower framework [24], which is an open-source library designed to streamline federated learning tasks on a cluster of machines.
-Flower’s high-level API eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it already supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [25] and Federated SGD (FedSGD) [14].
-Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, as it allows seamless integration of local models while maintaining centralized coordination on the server.
-2.4.2 FednnU-Net
-Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real and simulated environments.
-Its modular structure enables seamless integration not only with the original nnU-Net but also with its customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup.
-2.4.3 Distributed Setup
+2 depicts an overview of the AsymFedAvg pipeline and illustrates the process of updating the heterogeneous architectures in each training round.
+Implementation
+Federated Learning Environment
+For the federated communication, we select the Flower framework [36], which is an open-source library designed to streamline FL tasks on a cluster of machines.
+Flower’s high-level application programming interface (API) eliminates the need for delving into complex and time-consuming processes such as network communication between nodes and model aggregation, as it supports multiple FL optimization algorithms such as Federated Averaging (FedAvg), FedProx [37] and Federated SGD (FedSGD) [15].
+Additionally, Flower’s client-server architecture is well-suited for our multi-center setup, facilitating seamless integration of local models while maintaining centralized coordination on the server.
+FednnU-Net Framework
+Our framework maintains full functionality of the original nnU-Net while offering a federated extension that allows experimentation in both real-world and simulated environments.
+Focusing on usability, its modular structure enables seamless integration not only with the original nnU-Net but also with respective customized versions, allowing researchers to adapt their traditional centralized training methods to a federated setup.
+Key features include:
+  • Custom Architecture Selection: Users can select from nnUNet’s built-in architecture presets, including configurations optimized for high performance (e.g., residual encoder variants), even if they require more memory or compute.
+  • Per-Client Resource Configuration: GPU memory constraints can be individually defined for each client, enabling heterogeneous resource allocation scenarios such as in AsymFedAvg.
+  • Training Hyperparameter Control: Users can configure learning rate schedules, loss functions, optimizer settings, and number of training epochs through the native nnUNetTrainer class.
+  • Support for Pretrained Weights: Initialization from pretrained weights is supported, allowing faster convergence or domain adaptation across clients.
+  • Simulated and Real-World Federation: The same codebase supports both simulated federated setups (single machine, multi-client emulation) and deployment across physically distributed clients.
+This design ensures both high flexibility for research experimentation and practicality for real-world deployment across diverse clinical environments.
+Distributed Setup
 Our experimental framework consists of 4 computational nodes with varying hardware and system specifications (Table 2), resembling real-world clinical scenarios in low-resource countries.
@@
 Table 2: Hardware specifications of the computational nodes.
-Node CPU (Cores/GHz)* GPU (Memory)** OS ***
-Node 1 i7-8700K (12/3.20) GTX 1080 (8GB) Ubuntu 22.04
-Node 2 i7-8700K (12/3.20) GTX 1080 (8GB) Windows 10
-Node 3 i7-9700K (8/3.60) RTX 2080S (8GB) Ubuntu 22.04
-Node 4 i7-9700 (8/3.00) RTX 2080S (8GB) Ubuntu 22.04
-*GPU: Graphics Processing Unit, **CPU: Central Processing Unit, ***OS: Operating system
-3.  Datasets
-We evaluate our proposed methods on three distinct clinical applications, using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts.
+Node CPU (Cores/GHz) GPU (Memory) OS§
+Node 1 Intel i7-8700K (12/3.20) NVIDIA GTX 1080 (8GB) Ubuntu 22.04
+Node 2 Intel i7-8700K (12/3.20) NVIDIA GTX 1080 (8GB) Windows 10
+Node 3 Intel i7-9700K (8/3.60) NVIDIA RTX 2080S (8GB) Ubuntu 22.04
+Node 4 Intel i7-9700 (8/3.00) NVIDIA RTX 2080S (8GB) Ubuntu 22.04
+CPU: Central Processing Unit.  GPU: Graphics Processing Unit.  §OS: Operating System.
+Datasets
+We evaluate our proposed methods on three distinct clinical applications (i.e., breast, cardiac, fetal), using datasets that serve as valuable testbeds due to their multi-institutional and heterogeneous cohorts.
 The following sections provide a detailed description of the datasets employed in each use case.
-(a) MAMA-MIA
-(b) M&M
-(c) AIMIX
-Figure 3: Example images with annotations from each data center used in the study.
-3.1.  Breast MRI
-3.1.1 MAMA-MIA
-The MAMA-MIA dataset [26] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking AI algorithms in breast cancer imaging.
+Figure 3: Example images with annotations from each data center used in this work. From the top: Breast MRI, Cardiac MRI, Fetal US
+Breast MRI
+MAMA-MIA
+The MAMA-MIA dataset [38] is the largest publicly available collection of breast dynamic contrast-enhanced MRI (DCE-MRI) cases with expert primary tumor segmentations, providing an essential resource for advancing and benchmarking artificial intelligence (AI) algorithms in breast cancer imaging.
 It comprises 1,506 cases acquired between 1995 and 2016 from four cohorts within The Cancer Imaging Archive (TCIA): ISPY-1 (n=171), ISPY-2 (n=980), Duke (n=291), and NACT (n=64).
@@
 Slice
-Thickness (mm)
+Thickness ($m$)
 Number of
@@
 64
-3.1.2 EuCanImage:
-Through the European project EuCanImage [^1] https://eucanimage.eu/, we access a non-public breast MRI collection comprising 574 samples from three different countries.
+EuCanImage
+Through the European project EuCanImage [39], we access a non-public breast MRI collection comprising 574 samples from three different countries.
 The EuCanImage dataset represents a multi-institutional effort to create standardized imaging biobanks for cancer research and serves here as an external test set to validate our methodology on data acquired under protocols that differ markedly from those in MAMA-MIA.
@@
 Slice
-Thickness (mm)
+Thickness ($m$)
 Pixel
-Spacing (mm)
+Spacing ($m$)
 Field
@@
 312
-3.2.  Cardiac MRI
-3.2.1 M&Ms
+Cardiac MRI
+M&Ms
 To further validate our approach, we leverage the Multi-center, Multi-vendor & Multi-disease (M&Ms) cardiac MRI datasets.
-As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [27];[28] by excluding the overlapping samples based on anonymized patient IDs provided by each institution.
+As a part of our study, we curate a merged version of datasets that were used in the M&Ms challenges [40, 41] by excluding the overlapping samples based on anonymized patient IDs provided by each institution.
 This curation allows to reach the full potential of this data collection without the risk of replicated samples.
-The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct institutions.
-We make this collection publicly available [^2] This data resource will be released upon the acceptance of this article..
+The final, accumulated set consists of 543 unique, short-axis cardiac MRI scans with expert annotations from 5 distinct international institutions.
+We make this collection publicly available.
 The M&Ms datasets are particularly valuable for validation due to their inherent heterogeneity (Table 5), they include images from four different vendors (Siemens, GE, Philips, Canon), various scanner models (from older systems like Symphony to newer ones like Vantage Orian), and different field strengths (1.5T and 3.0T).
@@
 Slice
-Thickness (mm)
+Thickness ($m$)
 Number of
@@
 Studies
-Hospital Vall d’Hebron Siemens Symphony TIM (63.18%) Avanto Fit (15.48%) Other (21.34%) 9.5 13 1.5 (97.91%) 3.0 (2.09%) 239
-Clinica Sagrada Familia Philips Achieva 10 13 1.5 98
-Universitätsklinikum Hamburg-Eppendorf Philips Achieva 10 11 1.5 51
-Hospital Universitari Dexeus GE Signa Excite (75.24%) Signa HDxt (23.81%) Signa Explorer (0.95%) 9.9 12 1.5 (88.57%) 3.0 (11.43%) 105
+Hospital Vall d’Hebron Siemens
+Symphony TIM (63.18%)
+Avanto Fit (15.48%)
+Other (21.34%)
+9.5 13
+1.5 (97.91%)
+3.0 (2.09%)
+239
+Clinica Sagrada
+Familia
+Philips Achieva 10 13 1.5 98
+Universitätsklinikum
+Hamburg-Eppendorf
+Philips Achieva 10 11 1.5 51
+Hospital Universitari Dexeus GE
+Signa Excite (75.24%)
+Signa HDxt (23.81%)
+Signa Explorer (0.95%)
+9.9 12
+1.5 (88.57%)
+3.0 (11.43%)
+105
 Hospital Sant Pau Canon Vantage Orian 10 13 1.5 50
-3.2.2 ACDC:
+ACDC
 The Automated Cardiac Diagnosis Challenge (ACDC) dataset consists of 150 cardiac short-axis MRI examinations acquired at the University Hospital of Dijon (France) [3].
@@
 It is used for external validation of models trained on M&Ms dataset.
-3.3.  Fetal Ultrasound
-3.3.1 AIMIX
-With a focus on resource-constrained environments, the AIMIX [^3] https://aimix-erc.eu/ dataset was developed to advance inclusive image analysis in prenatal care.
+Fetal Ultrasound
+AIMIX
+With a focus on resource-constrained environments, the AIMIX [42] dataset was developed to advance inclusive image analysis in prenatal care.
 This dataset addresses critical challenges, including the scarcity of African medical data, limited computational resources, and a lack of clinical expertise.
@@
 In this study, we focus on abdominal segmentation, a crucial step in prenatal ultrasound analysis.
-To enable the evaluation on this multi-center dataset, we provide a set of 130 standard-plane images with masks corresponding to the fetal abdomen [^4] This data resource will be released upon the acceptance of this article..
-Our data processing workflow involved detecting the optimal abdominal plane from the each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [29], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician.
+To enable the evaluation on this multi-center dataset, we provide a set of 308 standard-plane images with masks corresponding to the fetal abdomen.
+Our data processing workflow involved detecting the optimal abdominal plane from each blind-sweep and segmenting it using a 4-point ellipse-fitting algorithm [43], utilizing a semi-automatic software by a researcher with more than five years of experience and reviewed by an expert clinician.
 Table 6 lists the number of samples and the mean resolution for each center.
@@
 SierraLeone (External) 561x743 178
-3.3.2 ACOUSLIC-AI Sierra Leone:
-For the external validation, we utilize the publicly available samples from ACOUSLIC-AI [^5] https://acouslic-ai.grand-challenge.org/datasets/ Challenge dataset [30].
+ACOUSLIC-AI Sierra Leone
+For the external validation, we utilize the publicly available samples from ACOUSLIC-AI Challenge dataset [44].
 We use a collection of 178 cases collected in Sierra Leone from three Public Health Units (PHUs).
@@
 This intermediate imaging quality makes the dataset an ideal candidate for assessing model generalizability across varying image qualities and clinical environments, ensuring robust validation of AI performance.
-4.  Experiments
-4.1.  Experimental setups
-The training process is conducted across four different setups:
-  • Single-center Training:  Each node independently trains its own nnU-Net model using only its local dataset.
+Experiments
+Experimental setups
+We adopt the vanilla nnU-Net architecture in both local and centralized training setups as the primary baselines for evaluating the performance of our proposed FednnU-Net framework.
+This choice is grounded in the fact that nnU-Net has established itself as a strong, standardized, and adaptive baseline for medical image segmentation tasks across a wide range of datasets [14].
+By using the same framework and memory constraints in local (per-institution) and centralized (pooled data) settings, we ensure a fair and controlled comparison that isolates the effect of the federation process itself, without introducing variability due to methodological, architectural or computational resource differences.
+This experimental setup allows for a clear evaluation of how well the federated setting bridges the gap between privacy preservation and segmentation accuracy, thus aligning with the core motivation of our study.
+To this end, the training process is conducted and empirically evaluated across four different setups:
+  • Local Training:  Each node independently trains its own nnU-Net model using only its local dataset.
 This setup serves as a baseline, reflecting the performance of a node/institution without participation in the federation.
-  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized model.
-This setup represents the theoretical upper bound of performance, though it is typically infeasible in real-world scenarios due to privacy constraints.
-  • Federated Fingerprint Training:   Utilizing our FFE approach (2), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node.
-As a result, each training site obtains identical network architecture during the nnU-Net’s self-configuration process.
-This approach approximates the nnU-Net’s behavior in the centralized training setup.
-  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, allowing nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
+  • Centralized Training:   All data is aggregated into a single dataset, which is used to train a centralized nnU-Net model.
+This setup represents the theoretical upper bound of performance, though it is typically infeasible and often not recommended in real-world scenarios due to significant privacy risks and data-sharing constraints.
+  • Federated Fingerprint Training:   Utilizing our FFE approach (see Methods), the local dataset fingerprints from all nodes are aggregated at the central node into one federated fingerprint and distributed back to each node.
+As a result, each training site obtains an identical network architecture during the nnU-Net’s self-configuration process and enables the utilization of a standard FedAvg aggregation method.
+Moreover, this approach approximates nnU-Net’s behavior in the centralized training setup.
+  • Asymmetric Aggregation Training:   Implemented using our AsymFedAvg strategy, this allows nodes to maintain their original nnU-Net configured architectures while sharing compatible parts of their networks.
 For our experimentation, we maintain nnU-Net’s default configuration parameters as described in [8].
 We test the 2D U-Net architecture on breast MRI and fetal ultrasound data and the 3D architecture on a cardiac MRI use case, consistently employing the default training length of 1000 epochs and using the last model checkpoint for evaluation across all federated nodes.
-Following other recent works on medical image segmentation [7, 31, 32, 33], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95).
+In federated scenarios, all models are aggregated after each epoch and sent back to the clients.
+Following other recent works on medical image segmentation [8, 45, 46], we evaluate the models with Dice Similarity Coefficient (DSC) and 95% Hausdorff distance (HD95).
 For internal validation, we report results as the average of a 5-fold cross-validation, while for external testing, we follow the original nnU-Net approach by generating ensemble predictions from all five folds.
 Moreover, each experiment uses identical data splits—for example, the n-th fold in the centralized setup corresponds directly to the set of n-th folds from centers used in the local setup.
+To assess statistical significance between segmentation methods, we apply the two-tailed Wilcoxon signed-rank test on paired DSC scores with 5% significance threshold.
 By adhering to these protocols, we try to minimize the influence of other factors on the final results, ensuring that any observed performance variations are primarily due to our federated learning adaptations rather than modifications to the base framework.
-4.2.  Results
-Table 8: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
+Results
+Table 7: Quantitative results for multi-center MAMA-MIA and EuCanImage datasets. HD95 values reported in $m$. Best results for privacy-preserving methods are in bold.
 Method ISPY1 ISPY2 DUKE NACT external GUMED external KAUNO external HCB
 Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
-Single-center 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
-AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
-FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
-Centralized 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
-Table 9: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in mm. Best results for privacy-preserving methods are in bold.
+nnU-Net (local) 0.612 17.60 0.739 19.03 0.616 33.56 0.551 23.24 0.579 35.64 0.636 31.72 0.544 46.30
+FednnU-Net AsymFedAvg 0.634 16.69 0.739 18.15 0.621 28.70 0.620 21.79 0.557 51.43 0.637 32.43 0.551 49.10
+FednnU-Net FFE 0.644 16.20 0.742 17.48 0.653 39.54 0.660 17.49 0.582 28.83 0.645 33.52 0.565 40.41
+nnU-Net (centralized) 0.654 17.28 0.750 19.84 0.657 58.99 0.675 17.35 0.582 52.44 0.644 41.17 0.569 57.98
+Table 8: Quantitative results for multi-center M&Ms and ACDC datasets. HD95 values reported in $m$. Best results for privacy-preserving methods are in bold.
 Method Hospital Vall d’Hebron Clinica Sagrada Familia Universitätsklinikum Hamburg-Eppendorf Hospital Dexeus Hospital Sant Pau external ACDC
 Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
-Single-center 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
-AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
-FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
-Centralized 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
+nnU-Net (local) 0.908 4.27 0.912 3.62 0.902 6.56 0.896 4.31 0.900 4.34 0.887 4.79
+FednnU-Net AsymFedAvg 0.909 4.20 0.913 3.70 0.907 4.23 0.897 4.29 0.902 4.07 0.889 4.58
+FednnU-Net FFE 0.911 4.05 0.914 3.67 0.922 3.30 0.904 3.98 0.904 3.87 0.897 3.87
+nnU-Net (centralized) 0.910 4.20 0.914 3.53 0.923 3.26 0.905 3.87 0.907 3.81 0.898 3.89
 Table 9: Quantitative results for multi-center AIMIX and ACOUSLIC-AI datasets.
@@
 Best results for privacy-preserving methods are in bold.
-Method Algeria Egypt Malawi Spain external Sierra Leone
+Method Algeria Egypt Malawi Spain
+external
+Sierra Leone
 Metrics DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$ DSC $\uparrow$ HD95 $\downarrow$
-Single-center 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
-AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
-FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
-Centralized 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
-(a) MAMA-MIA cross-center validation matrix
-(b) M&Ms cross-center validation matrix
-(c) AIMIX cross-center validation matrix
+nnU-Net (local) 0.952 11.87 0.931 24.21 0.926 26.83 0.942 29.82 0.831 53.20
+FednnU-Net AsymFedAvg 0.951 11.16 0.941 18.29 0.939 16.20 0.944 25.42 0.851 43.44
+FednnU-Net FFE 0.955 10.23 0.954 14.56 0.943 15.64 0.938 47.58 0.809 46.57
+nnU-Net (centralized) 0.954 10.23 0.953 14.50 0.943 15.55 0.932 55.76 0.804 54.34
+Breast MRI
+Table 7 summarizes segmentation performance on the MAMA-MIA dataset.
+Both FFE and AsymFedAvg outperform local training across all centers in DSC and HD95.
+FFE achieves average DSC improvements over local models on ISPY1 (3.2%, p=0.009), ISPY2 (0.3%, p=0.858), DUKE (3.7%, p=0.006), and NACT (10.9%, p¡0.001), also closely matching centralized results in DSC and surpassing them in HD95.
+Notably, AsymFedAvg yields the best HD95 on DUKE, reflecting its advantage in handling heterogeneous data (DUKE’s pixel size is 2× smaller; see Table 3).
+Fig.
+4 highlights this heterogeneity, with ISPY2 models generalizing better to DUKE due to similar resolutions.
+Sample predictions (Fig.
+5) show marked improvements for NACT and DUKE using federated methods.
+External validation on EuCanImage confirms FFE’s generalization, achieving DSC comparable to centralized training.
+KAUNO center results are consistent with MAMA-MIA, while GUMED’s drop likely reflects annotation discrepancies which are illustrated in Fig.
+5.
+Cardiac MRI
+For M&Ms, federated methods outperform local models on both metrics, except for Clinica Sagrada Familia, where HD95 is slightly better locally.
+FFE consistently achieves the best DSC scores across internal and external datasets (combined significance using Fischer’s method: p¡0.001), while AsymFedAvg shows limited benefit—likely due to the dataset’s homogeneity, where unified architectures are more suitable.
+Cross-center results (Fig.
+4) support this, with models generalizing well across unseen centers.
+Importantly, FFE matches centralized performance despite additional privacy constraints.
+Sample predictions (Fig.
+5) show consistent results for central slices, with more variability at the apex (ACDC) and atria (Vall d’Hebron).
+Fetal Ultrasound
+In the fetal abdomen segmentation task, FFE and AsymFedAvg differ markedly (Table 9).
+Strong heterogeneity (capture devices, resolution; Table 6) impairs generalization from Algeria, Egypt, and Malawi to higher-resolution Spanish samples (Fig.
+4).
+Unified architectures (FFE) are less optimal here: allowing Spain to retain its native architecture (AsymFedAvg) improves performance for external centers like Sierra Leone, notably the DSC raised to 0.851 (p ¡ 0.001).
+Fig.
+5 shows that local models often misidentify abdomen regions, especially for Sierra Leone, whereas federated approaches improve delineation.
 Figure 4: Cross-center DSC scores for isolated model training predictions across multiple centers.
-4.2.1 Breast MRI
-Table 8 presents the segmentation performance for the breast lesion MAMA-MIA dataset.
-Both proposed methods, FFE and AsymFedAvg, outperform the local training of individual centers in terms of DSC and HD95 across all datasets.
-Specifically, FFE method achieves an average DSC improvement of 3.2% on ISPY1, 0.3% on ISPY2, 3.7% on DUKE, and a notable 10.9% on NACT.
-Additionally, FFE maintains DSC scores close to the centralized approach, while surpassing it in terms of HD95, effectively bridging the performance gap in a decentralized framework.
-Interestingly for DUKE, AsymFedAvg outperforms every other method in terms of HD95.
-Since the DUKE dataset comprises samples of substantially different spatial characteristics (pixel resolution approximately 2 times higher, see Table 4), the adaptive nature of the asymmetric aggregation approach allows this center to maintain its initial architecture, while benefiting from the other centers.
-In this case, the HD95 gets reduced in contrast to unified architecture methods (FFE and centralized setting).
-The cross-center validation matrix in Fig.
-4 further exposes the dataset’s heterogeneity, where both ISPY2 and NACT trained models struggle to generalize well for the DUKE dataset, with the only exception of ISPY2, which has image resolution much closer to DUKE.
-Sample predictions presented in Fig.
-5 show the strong improvement of segmentation performance for NACT dataset, which struggles to perform well with its locally trained nnU-Net model.
-For DUKE, HD95 improvement can be observed for AsymFedAvg model prediction.
-External validation results for the EuCanImage dataset show that federated training with the FFE method achieves DSC scores on par with the centralized model.
-While the segmentation performance for the KAUNO center is comparable to the results achieved within MAMA-MIA dataset, demonstrating good generalizability.
-The GUMED center exhibits a notable decrease in performance, which is most likely attributed to discrepancies in annotation protocols rather than the data itself.
-An example of such difference is illustrated in Fig.
-7, where fine-grained model predictions fail to align with coarse ground truth regions provided by experts.
-Figure 5: Sample predictions of each method for Breast Cancer datasets.
-4.2.2 Cardiac MRI
-For M&Ms, similarly to previous dataset, both proposed federated learning methods outperform the local, non-collaborative training setup across both tested metrics.
-The only exception being Clinica Sagrada Familia, for which the HD95 is lower in the local setup by a minimal margin.
-FFE method consistently achieves the highest scores in terms of DSC for both internal and external datasets, while the adaptive AsymFedAvg does not reflect substantial benefits from federated training.
-Most likely, it is caused by dominantly homogeneous nature of M&Ms dataset, in which all data centers share similar sample characteristics making a common architecture method more suitable.
-This effect can be observed as well in the cross-center evaluation of locally trained models (see Fig.
-4) where each center managed to provide acceptable results for samples from other, unseen institutions.
-Moreover, it is important to note that the FFE method, despite the additional privacy requirement manages to provide almost identical performance to the centralized training scenario.
-Sample predictions for cardiac MRI dataset can be observed in Fig.
-6.
-All tested methods provide consistent results for central slices while having more visible disparities in regions close to the heart’s apex (sample predictions from ACDC) and heart’s atria (sample predictions from Vall d’Hebron).
-Figure 6: Sample predictions of each method for cardiac MRI datasets.
-4.2.3 Fetal Ultrasound
-For fetal abdomen segmentation task, considerable differences can be observed between FFE and AsymFedAvg methods; see Table 10.
-Due to strong data heterogeneity, i.e.
-discrepancies in capture devices and image resolution between centers (Table 7), the models trained with data from Algeria, Egypt or Malawi do not generalize well for higher resolution samples from Spain; see Fig.
-4c.
-In this case, having a unified training strategy for all centers does not prove to be the optimal solution.
-Allowing the largest center, i.e.
-Spain, to keep its original architecture allows to get a more appropriate model to the characteristics of the external center, i.e.
-Sierra Leone, while still benefiting from the federated training and contribution of other centers.
-This effect can be observed on sample predictions presented in Fig.
-7.
-The single-center nnU-Net model often struggled to identify the right region for abdomen delineation, which can be especially observed in the external validation samples from Sierra Leone.
-Figure 7: Sample predictions of each method for AIMIX dataset.
-5.  Discussion
-Our experiments across three diverse imaging use cases - breast MRI (MAMA-MIA), cardiac MRI (M&Ms), and fetal ultrasound (AIMIX) - show that FednnU-Net can reliably match, or even surpass conventional centralized strategies.
-These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while respecting the complexities of real-world clinical collaborations, such as data privacy and varied institution-specific protocols.
-Moreover, our findings underscore the critical need for methods that effectively handle data heterogeneity during federated model training.
-Whilst other, federated state-of-the-art segmentation methods may achieve higher accuracy with highly customized solutions for specific tasks, the primary objective of this work is to provide an easy-to-use, privacy-preserving solution accessible to researchers with varying levels of expertise.
-In doing so, we prioritize the broader applicability and real-world viability of the framework over achieving the absolute highest segmentation performance.
-The increasing emphasis on privacy-preserving, decentralized settings is rapidly becoming the standard in modern AI solutions — especially within healthcare, where patient data confidentiality is paramount.
-FednnU-Net aligns with this trend by enabling collaborative training without compromising data privacy.
-However, several aspects require further investigation:
+Figure 5: Sample predictions of each method for datasets used in the study.
+Discussion
+Key Contributions and Performance
+This study introduces FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that combines privacy preservation with high performance through two key innovations: Federated Fingerprint Extraction for standardized architecture sharing, and AsymFedAvg for flexible aggregation across heterogeneous node-specific models.
+Our experiments across three diverse medical imaging clinical tasks in (i) breast MRI (MAMA-MIA), (ii) cardiac MRI (M&Ms), and (iii) fetal ultrasound (AIMIX) - show that FednnU-Net reliably matches or surpasses the performance of centralized strategies across settings and modalities.
+These results add to a growing body of evidence illustrating that distributed model optimization can achieve comparable accuracy while simultaneously fulfilling the requirements of complex real-world clinical collaborations, including those related to data privacy and deviations in institution-specific protocols.
+Towards Trustworthy AI in Healthcare
+[…]
+While some federated segmentation methods report high accuracy by tailoring solutions to narrow, task-specific settings, this work focuses on delivering an accessible, privacy-preserving framework that emphasizes ease of use and adaptability.
+By prioritizing broad applicability and real-world viability, the framework is positioned to support diverse research settings and facilitate widespread adoption.
+Indeed, the ongoing shift towards privacy-preserving, decentralized deep learning solutions in healthcare is critical for protecting patient confidentiality [24].
+In addition to this, the FUTURE-AI General 2 guideline states the necessity of implementing rigorous data protection throughout the AI lifecycle, including the use of privacy-enhancing technologies and secure data governance.
+In the context of FL with FednnU-Net, data protection is operationalized by ensuring that patient data never leaves its local institutions of provenance, reducing the risk of sensitive information leakage during collaborative model training.
+To this end, FednnU-Net addresses critical privacy and security concerns inherent in decentralized healthcare AI solutions, fostering trust among stakeholders and ensuring compliance with ethical and regulatory standards.
+While FednnU-Net addresses critical requirements for privacy and local adaptation in line with FUTURE-AI guidelines, several important aspects still warrant further investigation:
   1. 1.
-    Scalability and Client Diversity: Future studies should test FednnU-Net with a larger number of nodes to better mimic real-world deployments, where numerous and geographically dispersed institutions may challenge model convergence and communication.
+    Scalability and Client Diversity: Future work is to assess FednnU-Net across an even larger, geographically dispersed federation with diverse patient populations [47] to study convergence dynamics at international, multi-institutional scales.
   2. 2.
-    Heterogeneous Computational Resources: Given that clients often have different processing capabilities, it is crucial to evaluate FednnU-Net under varied hardware constraints.
-Although our framework already supports such a scenario, this dimension was not explored in this current study.
+    Heterogeneous Computational Resources: While our framework already supports diverse node capabilities, further systematic ablation studies under varied hardware conditions, including low-resource settings [48], are motivated towards showing the versatility for clinical adoption of the framework.
   3. 3.
-    Enhanced Privacy Mechanisms: Beyond our current privacy measures, integrating techniques such as differential privacy [34, 35] and secure multiparty computation (SMPC) could further reduce the risk of data leakage, even in aggregated updates.
-In summary, although more accurate medical segmentation techniques exist, FednnU-Net fills a vital gap by balancing segmentation performance with ease-of-use and privacy preservation.
-As federated learning continues to mature and becomes an increasingly standard practice in healthcare, addressing these limitations through further evaluation and enhancements will be essential for broader adoption and improved robustness in clinical applications.
-6.  Conclusion
-In this study, we present FednnU-Net, a federated learning framework for 2D and 3D medical image segmentation that bridges the gap between privacy-sensitive setups and high-performance model training.
-Our framework introduces two methods: Federated Fingerprint Extraction (FFE) and Asymetric Federated Averaging (AsymFedAvg) that unlock the potential of distributed model training under privacy constraints.
-FFE enables construction of one, universal model architecture akin to centralized training, ensuring robustness across data centers.
-Alternatively, AsymFedAvg allows tailored model customization for each center to better align with local data characteristics.
-Our results demonstrate that both FFE and AsymFedAvg offer distinct performance benefits depending on the multi-center dataset characteristics, either achieving performance close to the centralized setting or surpassing it, particularly in highly diverse datasets.
-On top of that, both approaches preserve nnU-Net’s hallmark self-configuring capabilities, and due to our modular open-source implementation, they allow researchers to seamlessly translate their nnU-Net-derived centralized segmentation methods to the federated setting.
-Furthermore, we open source our framework and release data enhancing two public dataset collections to allow reproducibility of results.
-We hope that our contributions provide a robust foundation for further development and evaluation of medical image segmentation models in privacy-constrained settings.
+    Enhanced Privacy Mechanisms: FL improves data privacy by design, but it still does not guarantee full security.
+Attacks such as model inversion [49] or membership inference [50] can still extract sensitive information from shared model updates.
+Integrating differential privacy [51, 52, 53, 54], synthetic data generation [55, 56, 57], and secure multiparty computation[58] could further reduce the risk of data leakage, even in aggregated updates.
+Conclusion
+FednnU-Net fills a critical gap by balancing segmentation performance, ease of use, and privacy preservation.
+As FL gains adoption in healthcare [24], our FednnU-Net framework is a key component in enabling the translation of high-performance deep learning into privacy-compliant, scalable clinical solutions.
+Our framework enables researchers and practitioners to easily integrate nnU-Net, along with other dedicated frameworks and innovations built upon it, such as for instance Bayesian uncertainty estimation [21], continual learning [18], and curriculum learning extensions [16], into FL setups, thereby broadening their applicability under privacy constraints and supporting the increasingly prevalent multi-center collaborations.
+Continued evaluation, targeted enhancements, and the combination with such innovations is to further strengthen the versatility and robustness of the framework, paving the way for widespread, trustworthy deployment across diverse clinical environments.
+Data availability
+Our framework’s code is publicly available at https://github.com/faildeny/FednnUNet.
+The MAMA-MIA dataset access information is available at https://github.com/LidiaGarrucho/MAMA-MIA.
+The access to the updated version of M&Ms1 The M&Ms2 datasets and annotations for fetal US will be provided through https://github.com/faildeny/FednnUNet upon the publication of this article.
+For any queries regarding the data used in this study, please contact the corresponding author.
+Funding information
+This work received funding from the European Union’s Horizon Europe research and innovation programme under Grant Agreement No.
+101057849 (DataTools4Heart).
+This work has been supported by the European Union’s research and innovation programmes: Horizon Europe under Grant Agreement No.
+101057699 (RadioVal) and Grant Agreement No.
+101044779 (AIMIX), Horizon 2020 under Grant Agreement No.
+952103 (EuCanImage).
+Author contributions statement
+The study was conceived by G.S.
+and designed by G.S., C.M.-I., and K.L.
+G.S.
+developed the methods, wrote the program code, and conducted the experimental work.
+F.A.
+and C.M.-I.
+contributed to code development and experimental work.
+F.A., D.A.K., R.O., and E.R.P.
+contributed to manuscript writing.
+G.S.
+led the writing of the manuscript.
+G.S., F.A., D.A.K., and R.O.
+revised the article.
+C.M.-I., F.A., L.G., D.A.K., and E.R.P.
+worked on data preparation.
+M.B., K.G., X.B., and P.J.
+provided the imaging data and annotations for this study.
+M.B., K.G., X.B., P.J., C.M.-I., E.R.P., O.D., K.K., and K.L.
+reviewed the manuscript.
+K.K.
+and K.L.
+supervised the study.
+K.L.
+provided funding for the study.
+All authors have read and approved the final manuscript.
+Additional information
+Competing interests The authors declare no competing interests.